• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Templon

    Log entry time 12:08:40 on September28,2000

    Entry number 48432

    keyword=VDC cabling

    Found two VDC cables on right arm, and one cable on left arm, which were swapped
    at the TDC input. Swapped meaning twisted so that the wire ordering was
    backwards (cable was in correct input section, but upside down). Fixed.
    Started cosmic run to see whether problems in VDC wire spectra (which
    initiated this check) are fixed. Preliminary analysis indicates not all
    problems are fixed. Have NOT checked cable parity at output of preamp/disc
    cards (yet).